OpenStudy (anonymous):

If the surface area of a sphere is 144π cm2, what is the radius

OpenStudy (anonymous):

Use the formula for the surface area of a sphere.

OpenStudy (anonymous):

\[S=4\pi r^2\]

OpenStudy (anonymous):

is it 12?

OpenStudy (anonymous):

No. You made a small calculation error.

OpenStudy (anonymous):

144π=4π(r^2) solve for r

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

r^2=144π/4π=36

OpenStudy (anonymous):

r=sqrt(36)=6 is that right?

OpenStudy (anonymous):

Please don't post full answers, @mukushla. It is against the Code of Conduct.

OpenStudy (anonymous):

@Limitless ok...

OpenStudy (anonymous):

@mukushla, thank you.